OpenStudy (anonymous):

7. Solve the equation 5/4x-4 - 6/x+1 = -1 Help Please

OpenStudy (anonymous):

Multiply through by common denominator: (4x-4)(x+1)

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

Know how to use the equation function. Give me the right representation of your equation, or better still, draw it. thx

OpenStudy (anonymous):

so end up with 5x+5/(4x+4)(x+1) - 24x+24/(4x+4)(x+1) = -1/(4x+4)(x+1) ?

OpenStudy (anonymous):

Is it this?\[\frac{ 5 }{ 4x-4 } - \frac{ 6 }{ x+1} = -1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 5 }{ 4x-4 } (4x+4)(x+1)- \frac{ 6 }{ x+1}(4x+4)(x+1) = -1(4x+4)(x+1)\]

OpenStudy (anonymous):

ok so

OpenStudy (anonymous):

what's next

OpenStudy (anonymous):

\[\frac{ 5 }{ 4x-4 } (4x-4)(x+1)- \frac{ 6 }{ x+1}(4x-4)(x+1) = -1(4x-4)(x+1)\]

OpenStudy (anonymous):

Sorry, previous had wrong sign

OpenStudy (anonymous):

So, the identical brackets cancel out to:\[5(x+1)-6(4x-4) = -1(4x^2-4)\]

OpenStudy (anonymous):

Can you do the rest?

OpenStudy (anonymous):

Let me help you rearrange, so you can solve for the two values of x\[4x^2-19x+21 = 0\]Two factors you need are 12 and 7. Good luck :)